# DynaSeg: A Deep Dynamic Fusion Method for Unsupervised Image Segmentation Incorporating Feature Similarity and Spatial Continuity

### v1

@@ III-B Feature Extractor Network
The Feature Extractor Network, functioning as the primary processing stage, is tasked with extracting high-level features from input images. This pivotal role lays the foundation for subsequent stages by unravelling intricate details and capturing nuanced patterns within the input images. In our extended model, we explore two alternatives for feature extraction:
III-B1 CNN-based Network
We employ a CNN-based network that comprises $M$ convolutional components. Each component integrates $2D$ convolution, $Relu$ activation, and batch normalization, collectively forming a robust feature extraction pipeline. The output is a $p$-dimensional feature map $x_{n}$.
III-B2 Pre-trained ResNet with Feature Pyramid Network (FPN) Decoder
To expand the spectrum of feature extraction, we introduce an alternative feature extractor network—pre-trained ResNet [45]. This integration responds to the need for richer representations in image segmentation.
To tailor the ResNet architecture for segmentation, we eliminate the Global Average Pooling (GAP) and the Fully Connected Layer (FC) typically used in classification tasks. Furthermore, to restore the original size of the image, we employ a decoder. In contrast to a standard decoder that merely upsamples the extracted feature map, we opt for a Feature Pyramid Network (FPN) decoder. FPN involves lateral connections to fuse features from different resolutions. Implementing FPN on top of the output of the Conv5_x layer outputs a $p$-dimensional feature map $x_{n}$, making the height resolution features semantically strong.
This strategic combination leverages ResNet’s depth and FPN’s multi-scale feature integration, enabling the model to capture details at various granularity levels. We aim to achieve more robust and semantically meaningful image segmentation, focusing on practical descriptions.
This architectural choice facilitates the comparison and utilization of strengths inherent in both the CNN-based network and the pre-trained ResNet with a decoder (FPN). This contribution significantly enhances the overall performance of image segmentation, showcasing the effectiveness of our novel approach.
III-C Dynamic Weighting Scheme
In this section, we delve into the details of the proposed Dynamic Weighting Scheme aimed at mitigating the sensitivity issue of the balancing parameter $\mu$ in our unsupervised image segmentation model.
III-C1 Loss Function
@@
To strike a balance between feature similarity and spatial continuity, two pivotal criteria for distinguishing pixel clusters, our methodology incorporates specific loss functions. The feature similarity loss, denoted as $L_{sim}$ (Equation (1)), plays a crucial role in ensuring that pixels with similar features receive identical labels. This loss quantifies the cross-entropy between the normalized response map $r^{\prime}_{n}$ and the corresponding cluster labels $c_{n}$. Its minimization facilitates the extraction of precise attributes, contributing significantly to segmentation accuracy.
@@
Conversely, the spatial continuity loss $L_{con}$ acts as a high-pass filter, ensuring that spatially continuous pixels receive the same label. Defined by Equation (6), $L_{con}$ as the Manhattan Distance $L1$ Norm of horizontal and vertical differences in the response map $r^{\prime}_{n}$. This mitigates the deficiencies caused by superpixels [44], efficiently removing excess labels due to complex patterns or textures and ensuring spatially continuous pixels share the same label.
@@
While the combination of feature similarity ($L_{sim}$) and spatial continuity ($L_{con}$) losses in the unified function $L$ yields reasonably accurate unsupervised segmentation results, the parameter $\mu$ plays a crucial role and can lead to varying outcomes. The sensitivity to $\mu$ is illustrated in Figure (2), underscoring the challenge of selecting an appropriate value. As can be seen, for $\mu=50$ and $\mu=100$, the segmentation is coarse, resulting in sky, buildings, and coastal regions. However, the image is further segmented with $\mu=1$ and $\mu=5$, where buildings are further segmented into glass buildings, concrete buildings, and different floors. Although the authors in [13] argue that the value of $\mu$ is proportional to the coarseness of segmentation, We see that the results are not consistent, e.g. the segmentation for $\mu=50$ appears coarser than $\mu=100$. This poses a problem in practice. Such high sensitivity to the parameter means that for each dataset, this parameter has to be tuned extensively to obtain a result that is semantically more meaningful.
@@
TABLE I: Comparison of $mIOU$ for unsupervised segmentation on BSD500 and PASCAL VOC2012. Best scores are in bold.
@@
III-C2 Silhouette Score Phase
@@ III-D The Clustering Mechanism
In summary, our methodology presents a comprehensive approach to unsupervised image segmentation. The interplay of the Feature Extractor Network, Dynamic Weighting Scheme, and Clustering Mechanism, coupled with the innovative Silhouette Score Phase, contributes to state-of-the-art segmentation performance.
IV Experimental Results
@@ IV-A Experiment Setup
TABLE II: Comparison of different methods based on their mean Intersection over Union (mIoU) for ’Stuff’ categories and pixel accuracy (pAcc) on COCO-Stuff.
@@
IIC ResNet18 27.7 21.8
Picie ResNet18 31.48 50.0
DenseSiam ResNet18 24.5
HSG (Ke et al., 2022) ResNet50 23.8 57.6
ReCo+ (Shin et al., 2022) DeiT-B/8 32.6 54.1
DINO (Caron et al., 2021) ViT-B/8 41.9 74.9
@@
TABLE III: Comparison of mIOU for unsupervised segmentation on COCO-All.
@@
Modified Deep Clustering - 9.8
IIC ResNet18 6.7
Picie ResNet18 14.4
DenseSiam ResNet18 16.4
@@
IV-A1 COCO-Stuff
IV-A2 BSD500
IV-A3 PASCAL VOC2012
@@
Figure 3: Qualitative Results on select BSD500 and PASCAL VOC2012 images. Same color corresponds to the pixels being assigned the same clustering label by the algorithm. Please read Section IV-D for discussion on these results.
IV-B Evaluation
@@ IV-C Quantitative Results
IV-C1 COCO-All Dataset
Table III presents the results on the COCO-All dataset, comparing mean Intersection over Union (mIoU) scores. The two proposed DynaSeg versions, Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF), with different backbones, exhibit varying performance. Notably, SCF with ResNet-18 FPN achieves a state-of-the-art mIoU of 30.52, surpassing existing benchmarks. SCF outperforms DenseSiam by 14.12 and Picie by 16.08, establishing its leadership in unsupervised semantic segmentation.
IV-C2 COCO-Stuff Dataset
Table II summarizes the results for the COCO-stuff dataset, considering different backbones and architectures. SCF with CNN-Based achieves an mIoU Stuff of 42.41 and pAcc of 76.7, demonstrating its effectiveness. However, FSF with ResNet-18 FPN emerges as the new state-of-the-art, surpassing all methods with the highest mIoU Stuff of 54.10 and pAcc of 81.1.
IV-C3 BSD500 and PASCAL VOC2012 Datasets
Quantitative results on BSD500 and PASCAL VOC2012 datasets are detailed in Table I. SCF achieves a remarkable mIOU of 0.396 on PASCAL VOC2012, demonstrating its effectiveness. FSF consistently outperforms other methods on BSD500, emphasizing its robustness across various ground truth types. In addition to quantitative evaluations, we provide a demo [51] of incremental segmentation using DynaSeg on a sample image from the Pascal VOC2012 dataset. This demo is available in the supplemental video, showcasing our method’s application to real-world images.
@@ IV-D Qualitative Results
Figure 5: Qualitative Results on select Icoseg and Pixabay images. Same color corresponds to the pixels being assigned the same clustering label by the algorithm. Please read Section IV-D for discussion on these results.
@@
V Ablation Study
This section presents supplementary results examining the influence of the Silhouette Score Phase and the Feature Extractor through ablation on the COCO dataset. Additionally, we provide a comparison between FSF and SCF to further analyze their respective contributions.
@@ V Ablation Study
V-A Silhouette Score Impact
To evaluate the impact of integrating the Silhouette Score into our proposed dynamic weighting scheme, we conducted an ablation study, comparing segmentation results with and without the inclusion of Silhouette Score. Table IV presents the comparative results across different configurations.
@@ V-A Silhouette Score Impact
TABLE IV: Silhouette Score Impact
Framework Backbone Threshold mIoU
@@
V-B Effect of Feature Extractor
To analyze the impact of the choice of feature extractor, we compare the performance of CNN-Based and ResNet-18 with FPN in the FSF and SCF methods. Table V comprehensively compares segmentation results across different combinations of frameworks and backbones.
TABLE V: Comparison of framework and backbone variations.
@@ V-B Effect of Feature Extractor
V-C Comparison Between FSF and SCF
To gain insights into the differences between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we conduct a comparative analysis using Table V.
@@ V-C Comparison Between FSF and SCF
To further refine the comparative analysis between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we consider additional mIOU scores on BSD500 and PASCAL VOC2012 datasets, as presented in Table I.
@@
Comparing the two methods, it is evident that both FSF and SCF show sensitivity to the choice of the backbone architecture, with ResNet-18 consistently outperforming CNN-Based backbones. FSF, however, exhibits higher mIoU values for ”Stuff” across both backbones, emphasizing its effectiveness in capturing detailed and intricate features.
In conclusion, the choice between FSF and SCF should consider the specific requirements of the segmentation task, with FSF showing promise for scenarios involving complex and diverse categories, particularly when coupled with a ResNet-18 backbone.
VI Conclusion
In conclusion, our study introduces a state-of-the-art unsupervised image segmentation approach that effectively overcomes the limitations of supervised methods, which heavily depend on costly pixel-level annotations. The dynamic weighting scheme, Silhouette Score Phase, integration of a pre-trained ResNet feature extraction, and Feature Pyramid Network (FPN) decoding collectively contribute to superior performance. Remarkably, our method achieves the pinnacle of segmentation accuracy without extensive parameter tuning, showcasing adaptability across diverse datasets. Our approach establishes itself as a state-of-the-art solution through extensive quantitative evaluation, demonstrating its efficacy in real-world applications.
[Uncaptioned image] Boujemaa Guermazi is a researcher and academician in Electrical and Computer Engineering at Toronto Metropolitan University. His research interests encompass deep learning, image segmentation, and unsupervised machine learning. He earned his M.Sc Degree in Signals, Systems, and Data from the National Engineering School of Tunis in 2018. Before that, he completed a M.Sc Degree in Aircraft Operation Engineering - Pilot Engineer at the Aviation School of Borj El Amri in 2010. Boujemaa is dedicated to advancing the understanding and application of cutting-edge technologies in the engineering sciences.
[Uncaptioned image] Riadh Ksantini received the M.Sc. and Ph.D. degrees in Computer Science from the Université de Sherbrooke, Sherbrooke, QC, Canada, in 2003 and 2007, respectively. Presently, he is Associate Professor at the Department of Computer Science, College of IT, University of Bahrain, Adjunct Associate Professor at the School of Computer Science, within the Faculty of Science of the University of Windsor, and Adjunct Professor at the Department of Computer Science, Université du Québec à Montréal (UQAM). His research interests include Artificial Intelligence, Machine/Deep Learning, Pattern Recognition and Computer Vision.
[Uncaptioned image] Naimul Khan (Senior member, IEEE) is an associate professor in the electrical, computer, and biomedical engineering department at Toronto Metropolitan University. His research interest is in the user-centric intelligent systems combining machine learning and image processing. He has over 80 publications in this areas, backed by significant funding from government agencies and several awards.

### v3

@@ 3.2 Feature Extractor Network
The Feature Extractor Network, functioning as the primary processing stage, is tasked with extracting high-level features from input images. This pivotal role lays the foundation for subsequent stages by unravelling intricate details and capturing nuanced patterns within the input images. In our extended model, we explore two alternatives for feature extraction: a CNN-based Network comprising $M$ convolutional components and a pre-trained ResNet [49] with a Feature Pyramid Network (FPN) decoder. This exploration allows us to compare and leverage the strengths of each approach for improved image segmentation, ensuring robustness and efficiency in feature extraction.
3.2.1 CNN-based Network
Our CNN-based network consists of $M$ convolutional components, each integrating 2D convolution, ReLU activation, and batch normalization. This architecture forms a robust feature extraction pipeline, producing a $p$-dimensional feature map $x_{n}$. The network’s simplicity ensures computational efficiency, making it suitable for practical applications without the need for extensive computational resources.
Pooling layers often reduce the spatial resolution of feature maps, leading to the loss of fine-grained details crucial for accurate image segmentation. By avoiding pooling, our network retains the original resolution of the feature maps, preserving critical spatial information. This design choice enhances the network’s ability to produce high-quality, semantically meaningful segments without sacrificing computational efficiency. Additionally, the CNN-based approach’s robustness in handling high-dimensional data makes it particularly effective for various image segmentation tasks.
3.2.2 Pre-trained ResNet with Feature Pyramid Network (FPN) Decoder
To enhance our network’s feature extraction capabilities, we incorporate a pre-trained ResNet-18 model. This approach addresses the need for more prosperous and diverse image segmentation representations, leveraging the strengths of both ResNet and FPN.
  • ResNet Adaptation: We modify the ResNet-18 architecture by removing the Global Average Pooling (GAP) and Fully Connected (FC) layers, typically used for classification tasks. Instead, we replace the final fully connected layer with a convolutional layer to produce spatially preserved feature maps suitable for segmentation tasks. ResNet’s use of residual connections effectively mitigates the vanishing gradient problem, allowing for deeper network training and more accurate feature extraction.
  • FPN Decoder: We integrate a Feature Pyramid Network (FPN) decoder instead of employing a standard decoder. The FPN utilizes lateral connections to fuse features from different resolutions, enhancing the representation of multi-scale features. By applying the FPN on top of the high-level features from the ResNet (specifically the Conv5_x layer), we generate a $p$-dimensional feature map $x_{n}$. This approach retains high-resolution features that are semantically rich and detailed.
This strategic combination leverages ResNet-18’s depth and robust feature extraction capabilities along with FPN’s multi-scale feature integration. As a result, our model can capture and utilize details at various granularity levels, significantly improving image segmentation performance. The residual connections in ResNet-18 help in mitigating the vanishing gradient problem, allowing for deeper and more accurate feature extraction. This complements the FPN’s ability to integrate features at multiple scales, providing a comprehensive representation that enhances segmentation accuracy.
The dual approach of using both a CNN-based network and a pre-trained ResNet with FPN decoder allows us to compare and leverage the strengths of each method. The CNN-based network offers robustness and computational efficiency, making it suitable for scenarios with limited computational resources. In contrast, the pre-trained ResNet with FPN decoder provides richer and more diverse feature representations, which are crucial for capturing fine-grained details in complex images. By incorporating both approaches, we achieve a balanced and highly effective feature extraction mechanism that enhances the overall performance of our segmentation framework.
3.3 Dynamic Weighting Scheme
The Dynamic Weighting Scheme is a novel contribution of our method, designed to dynamically adjust the weighting parameter ($\mu$) during training iterations. This adaptive approach offers several key advantages:
- Flexibility and Robustness: Traditional methods often rely on fixed weighting parameters, which may not be optimal for all datasets or image complexities. Our dynamic weighting scheme continuously adapts $\mu$, ensuring an optimal balance between feature similarity and spatial continuity throughout training. This reduces the need for extensive manual parameter tuning, enhancing the model’s robustness and usability across diverse applications.
- Improved Performance: By adjusting $\mu$ dynamically, the model can emphasize feature similarity or spatial continuity as needed during different training stages. This adaptability leads to improved segmentation performance, as the model fine-tunes itself to the specific characteristics of the data being processed.
The dynamic adjustment of $\mu$ addresses specific challenges such as the difficulty in maintaining a consistent balance between feature similarity and spatial continuity and the sensitivity of traditional models to fixed parameter settings. By continuously evaluating and updating $\mu$, our scheme ensures that the model remains adaptable and effective across different training states.
3.3.1 Loss Function
@@
Our methodology incorporates specific loss functions to strike a balance between feature similarity and spatial continuity—two pivotal criteria for distinguishing pixel clusters. The feature similarity loss, denoted as $L_{sim}$ (Equation (1)), plays a crucial role in ensuring that pixels with similar features receive identical labels. This loss quantifies the cross-entropy between the normalized response map $r^{\prime}_{n}$ and the corresponding cluster labels $c_{n}$. Minimizing this loss facilitates the extraction of precise attributes, significantly contributing to segmentation accuracy.
@@
Conversely, the spatial continuity loss $L_{con}$ acts as a high-pass filter, ensuring that spatially continuous pixels receive the same label. Defined by Equation (6), $L_{con}$ as the Manhattan Distance $L1$ Norm of horizontal and vertical differences in the response map $r^{\prime}_{n}$. This mitigates the deficiencies caused by superpixels [47], efficiently removing excess labels due to complex patterns or textures and ensuring spatially continuous pixels share the same label. This loss also promotes homogeneity within each cluster, enhancing segmentation coherence and reducing noise within clusters.
@@
While the combination of feature similarity ($L_{sim}$) and spatial continuity ($L_{con}$) losses in the unified function $L$ yields reasonably accurate unsupervised segmentation results, the parameter $\mu$ plays a crucial role and can lead to varying outcomes. The sensitivity to $\mu$ is illustrated in Figure (2), underscoring the challenge of selecting an appropriate value. As can be seen, for $\mu=50$ and $\mu=100$, the segmentation is coarse, resulting in sky, buildings, and coastal regions. However, the image is further segmented with $\mu=1$ and $\mu=5$, where buildings are further segmented into glass buildings, concrete buildings, and different floors. Although the authors in [17] argue that the value of $\mu$ is proportional to the coarseness of segmentation, We see that the results are not consistent, e.g. the segmentation for $\mu=50$ appears coarser than $\mu=100$. This inconsistency poses a practical problem. The high sensitivity to the parameter means that extensive tuning is required for each dataset to obtain semantically meaningful results.
@@
Table 1: Comparison of $mIOU$ for unsupervised segmentation on BSD500 and PASCAL VOC2012. Best scores are in bold.
@@
3.3.2 Silhouette Score Phase
@@ 3.4 The Clustering Mechanism
In summary, the Dynamic Weighting Scheme in DynaSeg significantly enhances segmentation accuracy and flexibility by integrating several innovative approaches. The dynamically weighted loss function allows the model to adapt to different datasets without manual parameter adjustments, simplifying implementation across various segmentation networks. Adaptive clustering adjusts the number of clusters dynamically during training, enhancing the model’s flexibility and robustness to accommodate the varying complexities of different datasets. Joint optimization seamlessly integrates feature extraction and clustering, with the backpropagation of clustering losses directly influencing the CNN, overcoming limitations associated with predetermined boundaries and data imbalance. The spatial continuity loss further improves segmentation accuracy by preserving high-frequency details and promoting homogeneity within each cluster, resulting in more coherent segmentations with reduced noise. DynaSeg’s design eliminates the need for predefined boundaries or superpixels, providing enhanced flexibility and adaptability, making the segmentation process robust to input data variations and effective across a wide range of scenarios. Overall, these integrated features lead to more accurate, coherent, and adaptable segmentation results.
4 Experimental Results
@@ 4.1 Experiment Setup
4.1.1 COCO-Stuff
4.1.2 BSD500
@@
IIC [16] ResNet18 27.7 21.8
Picie [44] ResNet18 31.48 50.0
DenseSiam [45] ResNet18 24.5 -
HSG [53] ResNet50 23.8 57.6
ReCo+ [54] DeiT-B/8 32.6 54.1
STEGO [55] ViT-B/16 23.7 52.5
DINO + ACSeg [56] ViT-B/8 16.4 -
DINO + HP [57] ViT-B/8 24.6 57.2
DINO + CAUSE-TR [58] ViT-B/8 41.9 74.9
@@
Table 2: Comparison of different methods based on their mean Intersection over Union (mIoU) for ’Stuff’ categories and pixel accuracy (pAcc) on COCO-Stuff.
@@
Modified DC [59] - 9.8
IIC [16] ResNet18 6.7
Picie [44] ResNet18 14.4
DenseSiam [45] ResNet18 16.4
@@
Table 3: Comparison of mIOU for unsupervised segmentation on COCO-All.
4.1.3 PASCAL VOC2012
@@
Figure 3: Qualitative Results on select BSD500 and PASCAL VOC2012 images. Same color corresponds to the pixels being assigned the same clustering label by the algorithm. Please read Section 4.4 for discussion on these results.
4.2 Evaluation
@@ 4.3 Quantitative Results
4.3.1 COCO-All Dataset
Table 3 presents the results on the COCO-All dataset, comparing mean Intersection over Union (mIoU) scores. Our proposed DynaSeg models, particularly the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) versions, exhibit significantly improved performance compared to existing methods. Notably, DynaSeg - SCF with ResNet-18 and FPN Extractor achieves the highest mIoU of 30.52, surpassing benchmarks like DenseSiam and Picie by substantial margins (14.12 and 16.08, respectively). This establishes our approach as the new state-of-the-art in unsupervised semantic segmentation.
4.3.2 COCO-Stuff Dataset
Table 2 summarizes the results for the COCO-stuff dataset, considering different backbones and architectures. SCF with CNN-Based achieves an mIoU Stuff of 42.41 and pAcc of 76.7, demonstrating its effectiveness. However, FSF with ResNet-18 FPN emerges as the new state-of-the-art, surpassing all methods with the highest mIoU Stuff of 54.10 and pAcc of 81.1. These results are particularly noteworthy because they surpass the performance of other models, including those utilizing Vision Transformer (ViT) architectures. For example, methods like STEGO [55] and DINO + CAUSE-TR [58], which are built on ViT-B/16 and ViT-B/8 backbones respectively, do not achieve the same level of performance as our DynaSeg models. Specifically, DINO + CAUSE-TR, which previously set a high benchmark with an mIoU Stuff of 41.9, is significantly outperformed by our DynaSeg - FSF model.
The exceptional performance of the ResNet-18 FPN integration in our model highlights its ability to enhance feature extraction through multi-resolution processing. This capability allows our model to capture details at different scales more effectively, which is crucial for achieving superior segmentation performance, especially on datasets with diverse object scales and complexities. This underscores the advantage of our approach in leveraging the strengths of both ResNet and FPN, even against advanced transformer-based models.
4.3.3 BSD500 and PASCAL VOC2012 Datasets
Quantitative results on BSD500 and PASCAL VOC2012 datasets are detailed in Table 1. SCF achieves a remarkable mIOU of 0.396 on PASCAL VOC2012, demonstrating its effectiveness. FSF consistently outperforms other methods on BSD500, emphasizing its robustness across various ground truth types. In addition to quantitative evaluations, we provide a demo [62] of incremental segmentation using DynaSeg on a sample image from the Pascal VOC2012 dataset. This demo is available in the supplemental video, showcasing our method’s application to real-world images.
@@ 4.4 Qualitative Results
Figure 5: Qualitative Results on select Icoseg and Pixabay images. Same color corresponds to the pixels being assigned the same clustering label by the algorithm. Please read Section 4.4 for discussion on these results.
@@
5 Ablation Study
This section presents supplementary results examining the influence of the Silhouette Score Phase and the Feature Extractor through ablation on the COCO dataset. Additionally, we provide a comparison between FSF and SCF to further analyze their respective contributions. Furthermore, the integration of ResNet with the Feature Pyramid Network (FPN) significantly enhances feature extraction, leading to superior segmentation performance. This contribution highlights the effectiveness of our approach, particularly in handling objects at different scales and complexities.
@@ 5 Ablation Study
5.1 Silhouette Score Impact
To evaluate the impact of integrating the Silhouette Score into our proposed dynamic weighting scheme, we conducted an ablation study, comparing segmentation results with and without the inclusion of Silhouette Score. Table 4 presents the comparative results across different configurations.
@@ 5.1 Silhouette Score Impact
Framework Backbone Thr mIoU
@@
5.2 Effect of Feature Extractor
To analyze the impact of the choice of feature extractor, we compare the performance of CNN-Based and ResNet-18 with FPN in the FSF and SCF methods. Table 5 comprehensively compares segmentation results across different combinations of frameworks and backbones.
@@ 5.2 Effect of Feature Extractor
Table 5: Comparison of framework and backbone variations.
5.3 Comparison Between FSF and SCF
To gain insights into the differences between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we conduct a comparative analysis using Table 5.
@@ 5.3 Comparison Between FSF and SCF
To further refine the comparative analysis between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we consider additional mIOU scores on BSD500 and PASCAL VOC2012 datasets, as presented in Table 1.
@@
Both FSF and SCF perform better with a ResNet-18 backbone compared to CNN-based backbones. FSF, in particular, achieves higher mIoU values for ”Stuff” across both backbones, emphasizing its effectiveness in capturing detailed and intricate features. Therefore, FSF shows promise for scenarios involving complex and diverse categories, especially when coupled with a ResNet-18 backbone.
6 Conclusion
Our study introduces a state-of-the-art unsupervised image segmentation approach. The dynamic weighting scheme, Silhouette Score Phase, and integration of a pre-trained ResNet feature extraction and Feature Pyramid Network (FPN) decoding collectively contribute to superior performance. The dynamic weighting scheme, in particular, enhances segmentation accuracy and flexibility by dynamically adjusting loss weights and cluster numbers during training, simplifying implementation and improving adaptability across diverse datasets. Additionally, the joint optimization framework and spatial continuity loss promote coherent and balanced segmentations by preserving high-frequency details and ensuring homogeneity within clusters. Our method achieves high segmentation accuracy without extensive parameter tuning, showcasing its adaptability across diverse datasets.